Hello Sir, How many 3 digit nos. are there in which exactly 1 digit is 6?? Thanks

Hello Sir,


How many 3 digit nos. are there in which exactly 1 digit is 6??


Thanks

Grade:10

1 Answers

Sunil Raikwar
9 years ago
If hundred place is 6 then remaining place can be fill by 9*9=81 ways( excluding 6)
If tens place is 6 then remaining place can be fill 8*9=72 ways
Similarly when unit place is remaining place can be fill 8*9=72 ways
So required numbers whose exactly one digit is six= 225
Thanks & Regards
Sunil Raikwar
askIITians faculty